# Detecting Change Points of Covariance Matrices in High Dimensions

### v1

@@ 2 Change point analysis by a sequential LRT process
defines an asymptotic level $\alpha$-test for the hypotheses of a change point in the sequence $\mathbf{\Sigma}_{1},\ldots,\mathbf{\Sigma}_{n}$. The quantile $q_{\alpha}$ can be found numerically, replacing the asymptotic ratio $y$ in (2.12) by $p/n$. Next, we study the consistency of this test, for which we require an additional assumption on the spectrum of the matrices $\mathbf{\Sigma}_{1}$ and $\mathbf{\Sigma}_{n}$ in (2.2) before and after the change point. This assumption will be discussed later in Remark 1.
  1. (A4)
    The matrices $\mathbf{\Sigma}_{1}$ and $\mathbf{\Sigma}_{n}$ in (2.2) are simultaneously diagonalizable, that is, there exists an orthogonal matrix $\mathbf{Q}\in\mathbb{R}^{p\times p}$ such that
    $\mathbf{Q}\mathbf{\Sigma}_{1}\mathbf{Q}^{\top}=\operatorname{diag}(\lambda_{1}(\mathbf{\Sigma}_{1}),\ldots,\lambda_{p}(\mathbf{\Sigma}_{1})),\quad\mathbf{Q}\mathbf{\Sigma}_{n}\mathbf{Q}^{\top}=\operatorname{diag}(\lambda_{1}(\mathbf{\Sigma}_{n}),\ldots,\lambda_{p}(\mathbf{\Sigma}_{n})),$ (2.14)
    where $\lambda_{1}(\mathbf{\Sigma}_{1})\geq\ldots\geq\lambda_{p}(\mathbf{\Sigma}_{1})$ and $\lambda_{1}(\mathbf{\Sigma}_{n})\geq\ldots\geq\lambda_{p}(\mathbf{\Sigma}_{n})$ denote the ordered eigenvalues of the matrices $\mathbf{\Sigma}_{1}$ and $\mathbf{\Sigma}_{n}$, respectively. Suppose that these eigenvalues satisfy (uniformly with respect to $n\in\mathbb{N}$)
    $0<\lambda_{p}(\mathbf{\Sigma}_{1})\leq\lambda_{1}(\mathbf{\Sigma}_{1})<\infty,\quad 0<\lambda_{p}(\mathbf{\Sigma}_{n})\leq\lambda_{1}(\mathbf{\Sigma}_{n})<\infty.$
    Moreover,
    $E_{n}(t^{*})\to-\infty,$ (2.15)
    where $t^{\star}$ denotes the change point in (2.2) and $E_{n}(t^{*})$ is defined in (2.10).
Our next result shows that under this additional assumption, the test defined in (2.13) is consistent. A proof of this result can be found in Section A.1.
Theorem 2.
Suppose that Assumption (A1) and (A2) with $\delta>0$ are satisfied. Then, we have under Assumption (A4)
$M_{n}^{\textnormal{cen}}\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}-\infty.$
The theoretical guarantees for the test (2.11) are summarized in the following corollary, which is an immediate consequence of Theorem 1 and 2.
Corollary 1.
Assume that Assumptions (A1) and (A2) are satisfied.
  1. (a)
    (Level Control) Suppose that Assumption (A3) holds and that the parameter $\delta$ in Assumption (A2) satisfies $\delta>4.$ Then, we have under $H_{0}$
    $\lim_{n\to\infty}\mathbb{P}(M_{n}^{\textnormal{cen}}<q_{\alpha})=\alpha.$
  2. (b)
    (Consistency) Suppose that Assumption (A4) holds. Then, we have
    $\lim_{n\to\infty}\mathbb{P}(M_{n}^{\textnormal{cen}}<q_{\alpha})=1.$
If $H_{0}$ is rejected by the test (2.13), it is natural to ask for the location of the change point. For this purpose, we propose the following estimator.
@@
To ensure the convergence of this change-point estimator, we need the following assumption on the empirical spectral distributions of the matrices $\mathbf{\Sigma}_{1}$ and $\mathbf{\Sigma}_{n}$ under the alternative $H_{1}$
  1. (A1)
    The measure
    $\frac{1}{p}\sum\limits_{i=1}^{p}\delta_{(\lambda_{i}(\mathbf{\Sigma}_{1}),\lambda_{i}(\mathbf{\Sigma}_{n}))}$
    converges weakly to some measure $G$.
Theorem 3.
Suppose that Assumption (A1) and (A2) with $\delta>0$ are satisfied and that Assumption (A4) and (A1) hold. Then, we have under the alternative
$\hat{\tau}^{\star}\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}t^{\star},\quad n\to\infty.$
The proof can found in Section A.2.
Remark 1.
(1) The assumption that $\mathbf{\Sigma}_{1}$ and $\mathbf{\Sigma}_{n}$ are simultaneously diagonalizable, or equivalently, that both matrices commute, is of technical nature for deriving theoretical guarantees of our method under the alternative. Notably, such an assumption appears in related problems in random matrix theory. For instance, Li, Li and Yao (2018) study linear spectral statistics of sample covariance matrices with different population covariances. To understand the joint fluctuations of these dependent statistics, the population covariance matrices are assumed to be jointly diagonalizable. Another example is the work Liu, Aue and Paul (2015), in which the authors present a MP-type result for linear time series, assuming that the coefficient matrices of the linear process are simultaneously diagonalizable. The general nature of these problems is similar in spirit to our problem of understanding the behavior of $\Lambda_{n,t}^{(\textnormal{cen})}$ under the alternative, consisting of highly dependent eigenvalue statistics involving the population covariances $\mathbf{\Sigma}_{1}$ and $\mathbf{\Sigma}_{n}$ before and after the change point. However, under the null hypothesis of no change point, it is important to notice that $\Lambda_{n,t}^{(\textnormal{cen})}$ is invariant under the choice of $\mathbf{\Sigma}_{1}$. Thus, Assumption (A4) is solely needed for the analysis under the alternative hypothesis. Interestingly, the numerical experiments in Section 3 indicate that the performances of the proposed test and change point estimator do not suffer if this assumption is violated (see Section 3).
(2) The condition (2.15) ensures that the sequential likelihood ratio test statistic is consistent under the alternative. In fact, by (2.9) (which holds under under Assumption (A4)), the term $E_{n}(t^{*})$ determines the power of the test (2.13). This condition is akin to conditions (1.7) and (1.11) in Chen and Jiang (2018), ensuring the consistency of likelihood ratio tests for other testing problems.
3 Finite-sample properties
3.1 Data-scientific aspects
The necessity of $t_{0}$: The parameter $t_{0}$ ensures the applicability of the likelihood-ratio principle and is determined by the user. Parameters of this type appear frequently in monitoring high-dimensional covariance structures (see, for example, Ryan and Killick, 2023; Dörnemann and Dette, 2023; Dörnemann and Paul, 2024). In fact, there is one-to-one correspondence between $t_{0}$ and the minimum segment length parameter $\ell$ in Ryan and Killick (2023), and thus $t_{0}$ underlies the same paradigm as $\ell$ outlined in the aforementioned work. On the one hand, small values of $t_{0}$ are likely to increase the type-I error. In such cases, the maximal statistic will be dominated by covariance estimates corresponding to potential change points $t$ close to $p/n$ (or, by symmetry, close to $1-p/n$) which admit large eigenvalues. On the other hand, in many applications, the user may want to avoid large values for $t_{0}$, as such choices shrinkage the localization interval for change-point candidates. Therefore, it is important to understand how small the tuning parameter $t_{0}$ can be chosen without effecting the performance of the proposed method. Regarding the selection of $t_{0}$, it should first be noted that the parameter is unitless and does not need to be adapted to the scale of the model. By the design of the test statistic, a necessary lower bound will be $t_{0}>p/n\vee(1-p/n).$ In our simulation study, we found that the testing method performs stable if $t_{0}>(p/n+0.05)\vee 0.2$. If the user is primarily interested in estimating the change point location, they may select $t_{0}$ closer to the critical threshold $p/n\vee(1-p/n)$.
Non-simultaneously diagonalizable matrices $\mathbf{\Sigma}_{1}$ and $\mathbf{\Sigma}_{n}$: As explained in Remark 1, the technical assumption (2.14) allows us to derive theoretical guarantees of the proposed methodology under a structural break in the covariance structure. In our numerical study, we found that the change point detection and estimation is robust even under a larger class of alternatives. For instance, in the right panels of Figure 1 below, we report results for non-normally distributed data and randomly generated covariance matrices under the alternative with prescribed eigenvalues, see model (3.2).
@@ 3.2 Numerical experiments for change-point detection
We first consider independent standard normal distributed entries ($x_{11}\sim\mathcal{N}(0,1)$) in the matrix $\mathbf{X}$ and
$\mathbf{\Sigma}_{1}=\mathbf{I},\quad\mathbf{\Sigma}_{n}=\operatorname{diag}(1,\ldots,1,\underbrace{\delta,\ldots,\delta}_{p/2}),\quad\delta\geq 1,~{}$ (3.1)
as covariance matrices before and after the change point, where the case $\delta=1$ corresponds to null hypothesis (2.1). The empirical rejection probabilities of the test (2.13) are displayed in left panels of Figure 1 for $(n,p)=(600,50)$ (first row) and $(600,80)$ (middle row) and $(800,100)$ (third row) and various values of $\delta$. We observe that the test keeps its nominal level well and that the power increases quickly with $\delta$. For the sake of comparison, we also display the empirical rejection probabilities of the test proposed in Ryan and Killick (2023). As stated by these authors, this test is conservative, and we observe a substantial improvement with respect to power by the new test (2.13), which takes the dependencies of the statistics $\Lambda_{n,t}^{\textnormal{cen}}$ for different values of $t$ into account.
Next, we investigate the robustness of the new method with respect to the assumption of simultaneously diagonalizable covariance matrices before and after the change point. For this purpose, we consider the case where the matrices $\mathbf{\Sigma}_{1}$ and $\mathbf{\Sigma}_{n}$ before and after the change point are randomly generated with a prescribed spectrum, that is
$\mathbf{\Sigma}_{1}=\mathbf{I},\quad\mathbf{\Sigma}_{n}=\mathbf{U}_{\delta}\operatorname{diag}(1,\ldots,1,\underbrace{\delta,\ldots,\delta}_{p/2})\mathbf{U}_{\delta}^{\top},\quad\delta\geq 1,$ (3.2)
where $\mathbf{U}_{0}=\mathbf{I}$, and $\mathbf{U}_{\delta}$ are independent random matrices uniformly distributed on the orthogonal group for $\delta>1,i=1,2$. The independent entries in the matrix $\mathbf{X}$ are generated from a (uniform) $\mathcal{U}(0,1)$-distribution. The corresponding results are displayed in the right panels of Figure 1. Comparing these results with the left panels, we observe that the approximation of the nominal levels in the two models (3.1) and (3.2) is comparable. Despite the fact that model (3.2) does not satisfy Assumption (A4), the new test admits a favorable performance under this alternative, and we even observe an increase in power compared to model (3.1). In all cases under consideration, the new test outperforms the conservative method proposed by Ryan and Killick (2023) in terms of level approximation and power increase.
@@
3.3 Numerical experiments for the change-point estimation
In Table 2, we compare the mean, standard deviation and mean squared error of the new estimator $\hat{\tau}^{\star}$ in (2.16) with the RK estimator for the different alternatives in model (3.1) (with $\mathcal{N}(0,1)$-distributed independent entries in the matrix $\mathbf{X}$), where $t^{\star}=0.5$, $(n,p)=(600,50)$ (top), $(n,p)=(600,80)$ (middle) and $(n,p)=(800,100)$ (bottom). Note that the dimension is of comparable magnitude to the sample size, and therefore, the AHHR estimator cannot be computed and is therefore not included in the comparison. For example, for a dimension $p=50$, one requires at least a sample size of $(p+1)p/2+1=1276$ to calculate this estimator (some results for the AHHR estimator can be found in Table 4). We observe from the upper part of Table 2 that the new estimator (2.16) outperforms the RK estimator in all three cases under consideration. The smaller mean squared error of the new estimator (2.16) is caused by both a smaller bias and variance. In particular, the RK estimator admits a significant bias for moderately strong signals $\delta\approx 1.5.$
In Table 3, we display the results of the two estimators for model (3.2) with uniformly distributed data. The results are similar to those presented in Table 2 for model (3.1). It is noteworthy that the new estimator performs well, even though the technical assumption (A4) is not satisfied. Again, our method outperforms the alternative RK approach in terms of smaller mean squared error.
We conclude this section with a small comparison of the two estimators $\hat{\tau}^{\star}$ and RK with the estimator proposed by Aue et al. (2009) (AHHR) in the model (3.1). For this purpose, we select $t^{\star}=0.4$, and display the characteristics of the three change point estimators in Table 4. Note that the AHHR estimator can only be computed if the sample size is at least $p(p+1)/2+1$ and for this reason, we consider the cases $(n,p)=(200,10)$ (top), $(n,p)=(200,15)$ (bottom). As the dimension is relatively small compared to the sample size, we choose $t_{0}=0.1$. We observe that, even in such cases, the estimator AHHR admits a significant bias resulting in a larger MSE compared to the other two methods. Interestingly, the bias of RK increases as the signal strength $\delta$ increases from moderately to large values. In contrast, the new estimator $\hat{\tau}^{\star}$ has decreasing bias and standard deviation as $\delta$ increases. Moreover, the new estimator always outperforms RK indicated by a smaller mean squared error, and AHHR in the case $(n,p)=(200,15).$ For $(n,p)=(200,10),$ we observe that the mean squared error of AHHR is smaller for weak signal strength $\delta.$ However, even for large $\delta$, this method admits a significant bias and is therefore outperformed by our method.
$\delta$ 1.1 1.2 1.3 1.4 1.5 1.6 1.7 1.8 1.9 2
@@ 4.1 Proof of Theorem 1
$\hat{\mathbf{\Sigma}}_{i:j}^{(n)}$ $=\hat{\mathbf{\Sigma}}_{i:j}=\frac{1}{j-i+1}\sum\limits_{k=i}^{j}\mathbf{y}_{j}\mathbf{y}_{j}^{\top},\quad 1\leq i\leq j\leq n,$
$\hat{\mathbf{\Sigma}}$ $=\hat{\mathbf{\Sigma}}_{1:n}.$
@@
$\Lambda_{n,t}=\frac{\big{|}\hat{\mathbf{\Sigma}}_{1:\lfloor nt\rfloor}\big{|}^{\frac{1}{2}\lfloor nt\rfloor}\big{|}\hat{\mathbf{\Sigma}}_{(\lfloor nt\rfloor+1):n}\big{|}^{\frac{1}{2}(n-\lfloor nt\rfloor)}}{\big{|}\hat{\mathbf{\Sigma}}\big{|}^{\frac{1}{2}n}},\quad t\in(0,1).$ (4.1)
@@
$\mathbf{\Lambda}_{n}=((2\log\Lambda_{n,t}-\mu_{n,t})/n)_{t\in[t_{0},1-t_{0}]},$
@@
In the following theorem, we provide the convergence of the finite-dimensional distributions of $(\mathbf{\Lambda}_{n})_{n\in\mathbb{N}}.$
@@
Suppose that assumptions (A1), (A2) for some $\delta>0$, and (A3) are satisfied, and that $\mathbb{E}[x_{11}]=0$. For $n\to\infty$ and all fixed $k\in\mathbb{N}$, $t_{1},\ldots,t_{k}\in[t_{0},1-t_{0}]$, we have under $H_{0}$
@@
The asymptotic tightness of $(\mathbf{\Lambda}_{n})_{n\in\mathbb{N}}$ is given in the next theorem.
@@
Suppose that Assumptions (A1), (A2) with $\delta>4$ and (A3) are satisfied, and that $\mathbb{E}[x_{11}]=0$. Then, the sequence $(\mathbf{\Lambda}_{n})_{n\in\mathbb{N}}$ is asymptotically tight in the space $\ell^{\infty}([t_{0},1-t_{0}]).$
Proofs of these statements can be found in Section 4.1.1 and 4.1.3, respectively. Then, the weak convergence of $(\mathbf{\Lambda}_{n})_{n\in\mathbb{N}}$ towards a Gaussian process follows from the convergence of the finite-dimensional distributions (Theorem 4) and the tightness result (Theorem 5).
@@
Suppose that assumptions (A1), (A2) with $\delta>4$, (A3) are satisfied, and that $\mathbb{E}[x_{11}]=0$. Then, we have under the null hypothesis $H_{0}$ of no change point
@@
    Theorem 1 and Corollary 2 continue the line of literature on substitution principles in random matrix theory. When considering the spectral statistics of $\hat{\mathbf{\Sigma}}$ and $\hat{\mathbf{\Sigma}}^{(\textnormal{cen})}$, it was found by Zheng, Bai and Yao (2015) that their asymptotic distributions are linked by a substitution principle. This results says that one needs to substitute the location parameter $c_{n}$ in the CLT for the linear spectral statistics of $\hat{\mathbf{\Sigma}}$ by $c_{n-1}$ to account for the centralization in $\hat{\mathbf{\Sigma}}^{\textnormal{cen}}.$ A similar result has been found by Yin, Zheng and Zou (2023) for the linear eigenvalue statistics of the sample correlation matrix. However, it is important to emphasize that the test statistic $\Lambda_{n,t}^{\textnormal{cen}}$ considered in this work is a functional of several strongly dependent eigenvalue statistics, and therefore these results are not applicable. In fact, the analysis of $\Lambda_{n,t}^{\textnormal{cen}}$ requires a careful study, accounting for its intricate structure. These challenges will be faced even when restricting our focus to the case of one-dimensional distributions of $(\Lambda_{n,t})_{t}$, let alone considering the process convergence.
@@
    For the process convergence of $(\mathbf{\Lambda}_{n})_{n\in\mathbb{N}}$ in the space of bounded functions, the stronger moments condition (A2) with $\delta>4$ is needed, whereas moments of order $4+\delta$ for some $\delta>0$ are sufficient for the convergence of the finite-dimensional distributions of $(\log\Lambda_{n,t})_{t\in[t_{0},1-t_{0}]}$.
@@
$\hat{\mathbf{\Sigma}}^{\textnormal{cen}}=\frac{n}{n-1}\hat{\mathbf{\Sigma}}-\overline{\mathbf{y}}\overline{\mathbf{y}}^{\top},$

### v2

@@ 2 Change point analysis by a sequential LRT process
defines an asymptotic level $\alpha$-test for the hypotheses of a change point in the sequence $\boldsymbol{\Sigma}_{1},\ldots,\boldsymbol{\Sigma}_{n}$. The quantile $q_{\alpha}$ can be found numerically, replacing the asymptotic ratio $y$ in (2.10) by $p/n$. Then, Theorem 1 implies that the level of the test (2.11) can be asymptotically controlled under $H_{0}$, that is,
$\lim_{n\to\infty}\mathbb{P}\left(M_{n}^{\textnormal{cen}}<q_{\alpha}\right)=\alpha.$
3 Finite-sample properties
The necessity of $t_{0}$. The parameter $t_{0}$ ensures the applicability of the likelihood-ratio principle and is determined by the user. Parameters of this type appear frequently in monitoring high-dimensional covariance structures (see, for example, Ryan and Killick, 2023; Dörnemann and Dette, 2024; Dörnemann and Paul, 2024). In fact, there is one-to-one correspondence between $t_{0}$ and the minimum segment length parameter $\ell$ in Ryan and Killick (2023), and thus $t_{0}$ underlies the same paradigm as $\ell$ outlined in the aforementioned work. On the one hand, small values of $t_{0}$ are likely to increase the type-I error. In such cases, the maximal statistic will be dominated by covariance estimates corresponding to potential change points $t$ close to $p/n$ (or, by symmetry, close to 1p/n1-p/n) which admit large eigenvalues. On the other hand, in many applications, the user may want to avoid large values for $t_{0}$, as such choices shrink the localization interval for change-point candidates. Therefore, it is important to understand how small the tuning parameter $t_{0}$ can be chosen without affecting the performance of the proposed method. Regarding the selection of $t_{0}$, it should first be noted that the parameter is unitless and does not need to be adapted to the scale of the model. By the design of the test statistic, a necessary lower bound will be $t_{0}>p/n\vee(1-p/n).$ In our simulation study, we found that the testing method is stable if $t_{0}>(p/n+0.05)\vee 0.2$. If the user is primarily interested in estimating the change point location, they may select $t_{0}$ closer to the critical threshold $p/n\vee(1-p/n)$.
Estimating the change-point location. If $H_{0}$ is rejected by the test (2.11), it is natural to ask for the location of the change point. For this purpose, we propose the following estimator:
@@ 3 Finite-sample properties
In Section 3.2, we investigate the numerical performance of $\hat{\tau}^{\star}$ and compare it to the estimators of Aue et al. (2009) and Ryan and Killick (2023).
@@ 3.1 Numerical experiments for change-point detection
Recall that we observe the data $\mathbf{y}_{i}=\boldsymbol{\Sigma}_{i}^{1/2}\mathbf{x}_{i}$ for $1\leq i\leq n$, where $\boldsymbol{\Sigma}_{i}^{1/2}$ and $\mathbf{x}_{i}$ are not directly observed. We first consider independent standard normal distributed entries ($x_{11}\sim\mathcal{N}(0,1)$) in the vectors $\mathbf{x}_{i}$ and
$\boldsymbol{\Sigma}_{1}=\mathbf{I},\quad\boldsymbol{\Sigma}_{n}=\operatorname{diag}(1,\ldots,1,\underbrace{\eta,\ldots,\eta}_{p/2}),\quad\eta\geq 1,$~ (3.2)
as the covariance matrices before and after the change point, where the case $\eta=1$ corresponds to null hypothesis (2.1). The empirical rejection probabilities of the test (2.11) are displayed in the left panels of Figure 1 for $(n,p)=(600,50)$ (first row) and $(600,80)$ (middle row) and $(800,100)$ (third row) and various values of $\eta$. We observe that the test keeps its nominal level well and that the power increases quickly with $\eta$. For the sake of comparison, we also display the empirical rejection probabilities of the test proposed in Ryan and Killick (2023). As stated by these authors, this test is conservative, and we observe a substantial improvement with respect to power by the new test (2.11), which takes the dependencies of the statistics $\Lambda_{n,t}^{\textnormal{cen}}$ for different values of $t$ into account.
Next, we consider an adaptation of (3.3), where the matrix $\boldsymbol{\Sigma}_{n}$ is randomly generated with a prescribed spectrum, that is
$\boldsymbol{\Sigma}_{1}=\mathbf{I},\quad\boldsymbol{\Sigma}_{n}=\mathbf{U}_{\eta}\operatorname{diag}(1,\ldots,1,\underbrace{\eta,\ldots,\eta}_{p/2})\mathbf{U}_{\eta}^{\top},\quad\eta\geq 1,$ (3.3)
where $\mathbf{U}_{0}=\mathbf{I}$, and $\mathbf{U}_{\eta}$ are independent random matrices uniformly distributed on the orthogonal group for $\eta>1$. The independent entries in the matrix $\mathbf{X}$ are generated from a (uniform) $\mathcal{U}(0,1)$-distribution. The corresponding results are displayed in the right panels of Figure 1. Comparing these results with the left panels, we observe that the approximation of the nominal levels in the two models (3.2) and (3.3) is comparable. The new test shows a favorable performance under both alternatives (3.2) and (3.3) Notably, we observe an increase in power for (3.3) compared to (3.2), as (3.3) involves changes in both eigenvalues and eigenvectors, whereas (3.2) has changes only in the eigenvalues. In all cases under consideration, the new test outperforms the conservative method proposed by Ryan and Killick (2023) in terms of level approximation and power increase.
@@
Next, we investigate the performance of our method in the case where the change only affects the eigenvectors, but not the eigenvalues of $\boldsymbol{\Sigma}_{1}$ and $\boldsymbol{\Sigma}_{n}$. We consider the covariance matrices
$\boldsymbol{\Sigma}_{1}=\mathbf{Q}_{1}\operatorname{diag}(\underbrace{2,\ldots,2}_{p/2},\underbrace{1,\ldots,1}_{p/2})\mathbf{Q}_{1}^{\top},\quad\boldsymbol{\Sigma}_{n}=\mathbf{Q}_{2}\operatorname{diag}(\underbrace{2,\ldots,2}_{p/2},\underbrace{1,\ldots,1}_{p/2})\mathbf{Q}_{2}^{\top},$ (3.4)
where $\mathbf{Q}_{1},\mathbf{Q}_{2}$ are independent random matrices uniformly distributed on the orthogonal group. Note that this scenario corresponds to the alternative $H_{1}$ with overwhelming probability. Moreover, $\boldsymbol{\Sigma}_{1}$ and $\boldsymbol{\Sigma}_{n}$ share the same eigenvalues, so the spectrum does not change. Under $H_{0},$ we set $\boldsymbol{\Sigma}_{1}=\boldsymbol{\Sigma}_{n}$, and $\boldsymbol{\Sigma}_{1}$ is generated as above in (3.4). In Table 2, we display the empirical rejection rates of our proposed test. We observe that the proposed test attains full power against such alternatives.
$(n,p)$ (600,80) (800,100)
$x_{11}\sim\mathcal{U}(0,1)$ $H_{0}$ 0.058 0.066
$H_{1}$ 1.000 1.000
$x_{11}\sim\mathcal{N}(0,1)$ $H_{0}$ 0.068 0.062
$H_{1}$ 1.000 1.000
Table 2: Empirical rejection rates of the new test (2.11) for different values of $(n,p)$ and distributions for $x_{11}$, where $t_{0}=0.2,~t^{\star}=0.5$
$\eta$ 1.1 1.2 1.3 1.4 1.5 1.6 1.7 1.8 1.9 2
@@
3.2 Numerical experiments for the change-point estimation
In Table 3, we compare the mean, standard deviation and mean squared error of the new estimator $\hat{\tau}^{\star}$ in (3.1) with the RK estimator for the different alternatives in model (3.2) (with $\mathcal{N}(0,1)$-distributed independent entries in the matrix $\mathbf{X}$), where $t^{\star}=0.5$, $(n,p)=(600,50)$ (top), $(n,p)=(600,80)$ (middle) and $(n,p)=(800,100)$ (bottom). Note that the dimension is of comparable magnitude to the sample size, and therefore, the AHHR estimator cannot be computed and is therefore not included in the comparison. For example, for a dimension $p=50$, one requires at least a sample size of $(p+1)p/2+1=1276$ to calculate this estimator (some results for the AHHR estimator can be found in Table 5). We observe from the upper part of Table 3 that the new estimator (3.1) outperforms the RK estimator in all three cases under consideration. The smaller mean squared error of the new estimator (3.1) is caused by both a smaller bias and variance. In particular, the RK estimator admits a significant bias for moderately strong signals $\eta\approx 1.5.$
In Table 4, we display the results of the two estimators for model (3.3) with uniformly distributed data. The results are similar to those presented in Table 3 for model (3.2). Again, our method outperforms the alternative RK approach in terms of smaller mean squared error.
We conclude this section with a small comparison of the two estimators $\hat{\tau}^{\star}$ and RK with the estimator proposed by Aue et al. (2009) (AHHR) in the model (3.2). For this purpose, we select $t^{\star}=0.4$, and display the characteristics of the three change point estimators in Table 5. Note that the AHHR estimator can only be computed if the sample size is at least $p(p+1)/2+1$ and for this reason, we consider the cases $(n,p)=(200,10)$ (top), $(n,p)=(200,15)$ (bottom). As the dimension is relatively small compared to the sample size, we choose $t_{0}=0.1$. We observe that, even in such cases, the estimator AHHR admits a significant bias resulting in a larger MSE compared to the other two methods. Interestingly, the bias of RK increases as the signal strength $\eta$ increases from moderately to large values. In contrast, the new estimator $\hat{\tau}^{\star}$ has decreasing bias and standard deviation as $\eta$ increases. Moreover, the new estimator always outperforms RK indicated by a smaller mean squared error, and AHHR in the case $(n,p)=(200,15).$ For $(n,p)=(200,10),$ we observe that the mean squared error of AHHR is smaller for weak signal strength $\eta.$ However, even for large $\eta$, this method admits a significant bias and is therefore outperformed by our method.
@@ 4.1 Proof of Theorem 1
$\hat{\boldsymbol{\Sigma}}_{i:j}^{(n)}$ $=\hat{\boldsymbol{\Sigma}}_{i:j}=\frac{1}{j-i+1}\sum\limits_{k=i}^{j}\mathbf{y}_{j}\mathbf{y}_{j}^{\top},\quad 1\leq i\leq j\leq n,$
$\hat{\boldsymbol{\Sigma}}$ $=\hat{\boldsymbol{\Sigma}}_{1:n}.$
@@
$\Lambda_{n,t}=\frac{\big|\hat{\boldsymbol{\Sigma}}_{1:\lfloor nt\rfloor}\big|^{\frac{1}{2}\lfloor nt\rfloor}\big|\hat{\boldsymbol{\Sigma}}_{(\lfloor nt\rfloor+1):n}\big|^{\frac{1}{2}(n-\lfloor nt\rfloor)}}{\big|\hat{\boldsymbol{\Sigma}}\big|^{\frac{1}{2}n}},\quad t\in(0,1).$ (4.1)
@@
$\boldsymbol{\Lambda}_{n}=((2\log\Lambda_{n,t}-\mu_{n,t})/n)_{t\in[t_{0},1-t_{0}]},$
@@
In the following theorem, we provide the convergence of the finite-dimensional distributions of $(\boldsymbol{\Lambda}_{n})_{n\in\mathbb{N}}.$
@@
Suppose that assumptions (A-1), (A-2) for some $\delta>0$, and (A-3) are satisfied, and that $\mathbb{E}[x_{11}]=0$. For $n\to\infty$ and all fixed $k\in\mathbb{N}$, $t_{1},\ldots,t_{k}\in[t_{0},1-t_{0}]$, we have under $H_{0}$
@@
The asymptotic tightness of $(\boldsymbol{\Lambda}_{n})_{n\in\mathbb{N}}$ is given in the next theorem.
@@
Suppose that Assumptions (A-1), (A-2) with $\delta>4$ and (A-3) are satisfied, and that $\mathbb{E}[x_{11}]=0$. Then, the sequence $(\boldsymbol{\Lambda}_{n})_{n\in\mathbb{N}}$ is asymptotically tight in the space $\ell^{\infty}([t_{0},1-t_{0}]).$
Proofs of these statements can be found in Section 4.1.1 and A.2, respectively. Then, the weak convergence of $(\boldsymbol{\Lambda}_{n})_{n\in\mathbb{N}}$ towards a Gaussian process follows from the convergence of the finite-dimensional distributions (Theorem 2) and the tightness result (Theorem 3).
@@
Suppose that assumptions (A-1), (A-2) with $\delta>4$, (A-3) are satisfied, and that $\mathbb{E}[x_{11}]=0$. Then, we have under the null hypothesis $H_{0}$ of no change point
@@
    Theorem 1 and Corollary 1 continue the line of literature on substitution principles in random matrix theory. When considering the spectral statistics of $\hat{\boldsymbol{\Sigma}}$ and $\hat{\boldsymbol{\Sigma}}^{(\textnormal{cen})}$, it was found by Zheng et al. (2015) that their asymptotic distributions are linked by a substitution principle. This results says that one needs to substitute the location parameter $c_{n}$ in the CLT for the linear spectral statistics of $\hat{\boldsymbol{\Sigma}}$ by $c_{n-1}$ to account for the centralization in $\hat{\boldsymbol{\Sigma}}^{\textnormal{cen}}.$ A similar result has been found by Yin et al. (2023) for the linear eigenvalue statistics of the sample correlation matrix. However, it is important to emphasize that the test statistic $\Lambda_{n,t}^{\textnormal{cen}}$ considered in this work is a functional of several strongly dependent eigenvalue statistics, and therefore these results are not applicable. In fact, the analysis of $\Lambda_{n,t}^{\textnormal{cen}}$ requires a careful study, accounting for its intricate structure. These challenges will be faced even when restricting our focus to the case of one-dimensional distributions of $(\Lambda_{n,t})_{t}$, let alone considering the process convergence.
@@
    For the process convergence of $(\boldsymbol{\Lambda}_{n})_{n\in\mathbb{N}}$ in the space of bounded functions, the stronger moments condition (A-2) with $\delta>4$ is needed, whereas moments of order $4+\delta$ for some $\delta>0$ are sufficient for the convergence of the finite-dimensional distributions of $(\log\Lambda_{n,t})_{t\in[t_{0},1-t_{0}]}$.
@@
$\hat{\boldsymbol{\Sigma}}^{\textnormal{cen}}=\frac{n}{n-1}\hat{\boldsymbol{\Sigma}}-\overline{\mathbf{y}}\overline{\mathbf{y}}^{\top},$